OpenStudy (anonymous):

How do I find the equation of the normal line, (y=mx+b) to the curve at the given point? For example f(x)=sec(xy) - 1=tan(x/y) at (0,-1). I've already taken the derivative and end up with some really big complicated equation. Then, when I plug in the coordinate points to figure out the slope of the line, I get an answer that isn't possible like 0=-1. Any ideas?

OpenStudy (anonymous):

Get the derivative first :)

OpenStudy (anonymous):

Already have, but thanks for the tip!

OpenStudy (mertsj):

What did you get for the derivative?

OpenStudy (anonymous):

Well, if you get it, it's \[\Large \sec(xy)\tan(xy)[xy'+y]=\sec^2\left[\frac{x}y\right]\left(\frac{y-xy'}{y^2}\right)\]

OpenStudy (anonymous):

I actually wouldn't want to isolate yet, first, replace all the x's with 0, and all the y's with -1, and only then to isolate y' ^.^

OpenStudy (anonymous):

But then again, never mind :)

OpenStudy (anonymous):

I have seen your point...

OpenStudy (anonymous):

yeah, i'm not sure where I'm going wrong. It's either 0=1, or 0=-1

OpenStudy (anonymous):

Then maybe there simply is no escaping isolating at the first expression?

OpenStudy (anonymous):

I don't understand...what's the original equation you're given?

OpenStudy (mertsj):

Maybe the tangent is vertical.

OpenStudy (anonymous):

If you do isolate, I reckon it's going to be a 1/0 slope, so that's vertical. Then its normal line must be horizontal, no?

OpenStudy (mertsj):

Draw the graph of the function and see if it looks like that tan is vertical at the given point. (use wolf)

OpenStudy (anonymous):

Better, use wolf to actually get the derivative :D

OpenStudy (anonymous):

how in the world did you get that isolated?

OpenStudy (anonymous):

I didn't. Wolfram did.

OpenStudy (anonymous):

But if plugging x = 0 and y = -1 yields a false equation, it must mean that y' does not exist at that point. So it must be a vertical slope :)

OpenStudy (anonymous):

I plugged them in and got y' = -1/0. Does that mean the normal line for that equation would be 0/1, making it a horizontal line? If that's true, then I can just plug in the coordinate point into y=0(x)+b, right?

OpenStudy (anonymous):

Yeah, that's right :) Now you have a point (0 , -1) and a slope m = 0 Now get the line equation ^.^

OpenStudy (anonymous):

I got y=-1.

OpenStudy (anonymous):

And there's your normal line :)

OpenStudy (anonymous):

Awesome! Thanks for all the help.

OpenStudy (anonymous):

Anytime... well not really ^.^

OpenStudy (anonymous):

I was gonna say.. I got some more problems if you wanna help

OpenStudy (anonymous):

Post them. I'll see what I can do ^.^

OpenStudy (anonymous):

I posted one other one, the rates of change one. It's closed now, but think you can help with that one?

OpenStudy (anonymous):

Saw it, it makes me dizzy, I think I might do more harm than good If I tackle that :)

OpenStudy (anonymous):

Darn, thanks though!

OpenStudy (anonymous):

Sorry :)